# PanoLAM: Large Avatar Model for Gaussian Full-Head Synthesis from One-shot Unposed Image

### v1

@@ 2 Related work
BP4D-S 41 front Ours-Front 48,135 $72^{\circ}$
Nersemble 222 front Ours-360 117,186 $360^{\circ}$
Figure 1: Overall Framework. Given an unposed head image as input, PanoLAM involves two branches to achieve single-pass 3D Gaussian head reconstruction: a point-based transformer for coarse-to-fine point shape reconstruction and point features extraction, and a spherical triplane transformer to distill prior knowledge from the pretrained 3D GAN. The unstructured point features and the queried and aggregated structured spherical triplane features are concatenated for high-fidelity Gaussian head regression. We add auxiliary triplane rendering in the triplane-branch for supervision during training.
3 Methodology
3.1 Overview
As shown in Fig. 1, given an unposed reference image, our framework generates the Gaussian full-head in an efficient coarse-to-fine and joint point-triplane manner, where dense point cloud and point features are densified from sparse ones and fused with the spherical triplane features for dense Gaussian head regression. As shown in the upper part of Fig. 1, in the coarse stage of the point branch, sparse vertices from the FLAME vertices interact with the extracted multi-level image features by stacked cross-attention modules for feature extraction and coarse Gaussian point reconstruction. The dense reconstruction stage leverages barycentric weights from FLAME subdivision to densify the coarse Gaussian point position and features for dense Gaussian points regression. To fully leverage the prior knowledge residing in the spherical triplane of 3D GAN, as shown in the bottom branch of Fig. 1, a spherical triplane branch is also introduced, and an efficient aggregation mechanism is proposed to aggregate the structured features from the spherical triplane using the densified point cloud, which are then concatenated with the unstructured point features for high-fidelity Gaussian head reconstruction.
3.2 Coarse-to-fine Gaussian Head Avatar Generation
The number of Gaussian points is one key factor that can affect the fidelity of the reconstructed Gaussian head, especially for the head avatar that contains sharp details like hair strands, mustaches, and wrinkles. The total number of Gaussian points should be large enough to produce these details. However, the large number of Gaussian points is hard to optimize during training, especially in our feed-forward pipeline, where the iterative densification operation in the traditional 3D Gaussian Splatting training pipeline is not applicable. To resolve this problem, unlike previous work [30] that directly regresses the dense point cloud, we introduce a coarse-to-fine mechanism. Specifically, sparse Gaussian points are first reconstructed for coarse shape reconstruction and then densified to dense Gaussians for high-fidelity details crafting. There are two advantages in the proposed design. Firstly, the sparse number of points enables efficient cross-attention with the image features for textures and shape features extraction, saving large computation and memory costs compared with a large number of points. Secondly, the sparse point can be well supervised and optimized for large deformation to carve the shape of the head, especially in regions that require large deformation from the original FLAME model, e.g., long hairs. Dense Gaussian points upsampled from these correctly deformed sparse points can focus more on texture details crafting, which are much easier to optimize.
Coarse Gaussian Head Generation
To fully leverage the head shape prior that resides in the FLAME model, we initialize the coarse Gaussian point with sparse vertices (about 5000 points) on the neutral FLAME model in the canonical space. Our target is to regress the deformation offset for shape refinement of regions that FLAME cannot model, e.g., long hairs, glasses, and caps. To supervise the shape deformation with RGB input images, Gaussian attributes for each point are also regressed for differentiable Gaussian Splatting.
To get started, we assign positional embedding to each point and utilize learnable multi-layer perceptron (MLP) modules to project the channel of features into the token channel $C_{t}$ of transformers as follows:
$F_{P_{0}}=MLP(\gamma(V)),$ (1)
where $V$ is the spatial position of each vertice and $\gamma$ the L_\_frequency sinusoidal encoding as in NeRF [54]. To extract the texture and shape information in the given unposed image, we utilize the pre-trained DinoV2 [57] for feature extraction. Inspired by  [64, 30], we fuse features derived from both shallow and deep layers to obtain both local and global image features $F_{I}$. Specifically, an MLP fuses the $\{5,12,18,24\}$ layers of features in DinoV2 into $C_{t}$ token channels of features as:
$F_{I}=MLP(\mathcal{C}(F_{D_{5}},F_{D_{12}},F_{D_{18}},F_{D_{24}}),$ (2)
Given mapped point features $F_{P}$ and extracted image features $F_{I}$, we utilize $L_{\mathcal{A}}$ layers of stacked cross-attention modules $\mathcal{A}=\{\mathcal{A}_{i}\}_{i=1}^{L_{\mathcal{A}}}$ from Transformer [75] for feature extraction from image to point cloud as follows:
$F_{P_{i}}=\mathcal{A}_{i}(F_{P_{i-1}},F_{I}),$ (3)
$\{c_{k},o_{k},s_{k},R_{k},O_{k}\}^{M_{C}}_{k=1}=\mathcal{D}(F_{P_{k}}),$ (4)
Efficient Point Cloud Densification
After the reconstruction of the coarse Gaussian head, we obtain a sparse point cloud that describes the shape of the target head avatar in the image, each point is also attached with rich features extracted from the image. We then densify these points and features for high-fidelity dense Gaussian head reconstruction. Unlike previous methods [44, 82] that require a network trained to upsample sparse points, we leverage the topology in the FLAME mesh and propose an efficient densification strategy that utilizes the barycentric weights from mesh subdivision to densify the point cloud and feature. Our key observation is that after the deformation in the coarse reconstruction stage, the topology between each point mostly remains as the original FLAME vertices. Thereafter, we can precompute the barycentric weights for point cloud densification utilizing mesh subdivision on FLAME. Specifically, given the original FLAME model, we perform subdivision for faces that are larger than a threshold $\lambda_{area}$, the position of newly added vertices $v^{\prime}$ can be computed from the three vertices of the original face with barycentric weights as:
$v^{\prime}=\sum_{i=1}^{3}({w_{i}\cdot v_{i}}),$ (5)
where $w_{i}$ and $v_{i}$ are the barycentric weight and vertex coordinate of the original face, respectively. After the deformation in the coarse reconstruction stage, each vertex moves to a new position w.r.t the predicted offset as $p_{i}=v_{i}+O_{i}$. Then new added point $p^{\prime}$ with its attached point features $p_{F}^{\prime}$ can be interpolated with barycentric weights as:
$p^{\prime}=\sum_{i=1}^{3}({w_{i}\cdot p_{i}}),$ (6)
$p_{F}^{\prime}=\sum_{i=1}^{3}({w_{i}\cdot p_{F_{i}}}),$ (7)
where $p_{F_{i}}$ is the extracted features from the last cross-attention blocks in Equ. (9). In this way, we can efficiently densify the point and point features.
Dense Gaussian Head Generation.
After getting the dense point position with dense point features, we can utilize them for dense Gaussian head regression. Rather than using only unstructured point features for prediction, we also aggregate features from the structured spherical triplane for more effective Gaussian attributes regression, which we will discuss in the next section. Denote $F_{P_{i}}$ the $i_{th}$ point feature and $F_{T_{i}}$ the aggregated spherical triplane feature for point $P_{i}$, the dense Gaussian head attributes and a small deformation residual can be regressed similarly to the coarse one as:
3.3 Spherical Triplane Prior Knowledge Distillation
In the previous section, we obtain the dense point cloud with its attached point features extracted from the unposed image. Inspired by TriplaneGaussian [96], where structured triplane features are efficiently sampled by the unstructured point cloud for better Gaussian attributes regression, we introduce a spherical triplane branch to distill the knowledge from the trained SphereHead [43]. However, we find that the point-based query strategy in TriplaneGaussian cannot fully fetch valuable features from the spherical triplane. Therefore, we analyze the inner rendering mechanism of spherical triplane and propose a novel ray-based sampling strategy to aggregate multiple layers of features effectively. Specifically, we initialize the $H_{T}\times W_{T}\times 6\times C_{t}$ spherical triplane features with learnable query features, which are then flattened into spherical triplane token features $F_{T}\in\mathbb{R}^{N_{T}\times C_{t}}$, where $N_{T}=H_{T}\times W_{T}\times 6$. These learnable features then interact with the image features $F_{I}$ with stacked cross-attention blocks $\mathcal{A_{T}}=\{\mathcal{A_{T}}_{i}\}_{i=1}^{L_{\mathcal{A_{T}}}}$ similar to the learnable point features as:
$F_{T_{i}}=\mathcal{A_{T}}_{i}(F_{T_{i-1}},F_{I}),$ (9)
where $F_{T_{i}}$ is the $i_{th}$ layer of spherical triplane features in the stacked cross-attention layers. The final output features are then projected by MLPs to the same channel as the spherical triplane in SphereHead. We then need to query the feature from the spherical triplane to enhance the point features. One vanilla way is to follow TriplaneGaussian and use our densified point cloud to fetch the corresponding feature by projection. However, we find that our point clouds are distributed on the shape surface similar to a mesh, and such a single-layer query strategy cannot fully aggregate the corresponding features of each point from the spherical triplane. That’s because the triplane designed for NeRF-based rendering requires ray marching to aggregate multiple points(/layers) of features for rendering. Using a single point cloud on the shape surface cannot fully fetch valuable features from the spherical triplane. While ray marching is computationally consuming and inefficient, we propose to sample multiple layers of points from the single-layer point cloud for feature aggregation. Specifically, we sample 4 virtual cameras that can capture the front, left, right, and back sides of the head. As is shown in Fig. 2, for each camera, we cast a ray from the camera center to each point $P_{i}$, then $K_{m}$ number of points $P^{\prime}=\{P^{\prime}_{i}\}_{i=1}^{K_{m}}$ are sampled along the ray near the point. Each point then queries features from the spherical triplane by projection. We then concatenate and aggregate these queried features with MLPs to obtain the final spherical triplane features as:
$F_{T_{i}}=MLP(\mathcal{C}(\{F^{\prime}_{T_{i}}\}_{i=1}^{K_{m}})),$ (10)
where $F^{\prime}_{T_{i}}$ is the spherical triplane feature queried by sampled point $P^{\prime}_{i}$ ; $\mathcal{C}$ is the concatenation operation; and $F_{T_{i}}$ is the final aggregated feature for point $P_{i}$. In this way, each point gets its aggregated spherical triplane feature. These aggregated spherical triplane features can be concatenated with the point feature for dense Gaussian attribute regression as in Formula (8).
To supervise this branch in distilling the knowledge from SphereHead, we add the frozen decoding header from SphereHead to decode the spherical triplane feature to the target image and supervise it with the ground truth image during training. Specifically, as shown in the bottom right part of Fig. 1, for each target view, we project the point feature to the image plane by the Gaussian rasterizer and then use the frozen decoder from SphereHead to decode the target image from the projected spherical triplane features.
3.4 Large-scale Training Data Generation
@@
The scale of the training dataset is also a key factor in enabling the generalizability of a trained model. However, obtaining a large dataset with diverse head assets is challenging. Though various datasets have been proposed for 3D head modeling, as shown in Table 1, they have limited diversity of subjects, which hinders the generalization capability of learning-based models. Inspired by previous methods [62, 8, 28, 78] in various fields that model trained from synthesis data can generalize to real-world scenarios during inference, we leverage pretrained priors from 3D GANs and generate a large-scale, diverse dataset for training. Specifically, we leverage two prior models for data generation, EG3D for multiview front face images synthesis, and SphereHead for 360 view images generation. Our observation is that the EG3D generates a different distribution of images compared to SphereHead (e.g., diverse hairstyles and caps appeared in samples), while SphereHead is able to generate 360 images. The combination of two model spaces enlarges the diversity of the generated images and empowers better generalizability for learning-based models. However, bad cases may occur in the generated dataset, which can hurt the model. Therefore, we remove bad cases from the sampled images manually, ending up with 48,135 subjects in Ours-Front and 117,186 subjects in Ours-360 datasets.
@@ 3.5 Optimization and Regularization
In the training phase, we randomly select $N_{v}$ different views of images rendered from the same triplane. We randomly choose one as the reference image for the Gaussian full-head reconstruction, while the others serve as novel view images for prediction and training. Note that the two branches in our framework both render images for supervision. We ensure the accuracy of the rendered RGB images from each branch by comparing them with the ground truth target images, utilizing a combination of $\mathcal{L}_{1}$ loss and perceptual loss for supervision.
@@
Additionally, for the sparse and dense Gaussian Splatting branch, we render the silhouette and supervise it using $\mathcal{L}_{1}$ loss, referred to as $\mathcal{L}_{mask}$. And for the spherical triplane branch, we supervise the extracted triplane feature with the triplane feature in the pretrained SphereHead with $\mathcal{L}_{1}$ loss, denoted as $\mathcal{L}_{tri}$. The total loss function is a weighted sum as:
$\mathcal{L}=\lambda_{3}\mathcal{L}_{rgb}^{G}+\lambda_{4}\mathcal{L}_{rgb}^{T}+\lambda_{5}\mathcal{L}_{mask}+\lambda_{6}\mathcal{L}_{tri},$ (12)
where $\mathcal{L}_{rgb}^{G}$ denotes the RGB image loss for both sparse and dense Gaussian Splatting and $\mathcal{L}_{rgb}^{T}$ for the auxiliary triplane branch, both denoted as in Formula (11).
@@ 4.1 Experiments Setting
Our framework is implemented using PyTorch, with the DINOv2 image feature extraction backbone frozen. The Transformer component comprises 10 layers of basic blocks, featuring 16 attention heads and a feature dimension of 1024 ($C_{t}=1024$). The extracted features are translated into Gaussian attributes through three multilayer perceptron (MLP) layers. The training of the network is conducted over 100 epochs using the Adam optimizer and a cosine warm-up scheduler. Various hyperparameters are empirically set to optimize performance, including $N_{v}=6$, $\lambda_{1}=\lambda_{2}=1.0$, $\lambda_{3}=1.0$, $\lambda_{4}=0.1$, $\lambda_{5}=1.0$, $\lambda_{6}=0.0001$, $\epsilon_{O_{C}}=0.15$, $\epsilon_{O_{D}}=0.056$, $M_{D}=10K$, $K_{m}=32$, $\lambda_{area}=2.0e-6$, $H_{T}=W_{T}=64$.
@@
For our training and testing, we employ the large-scale dataset that we have generated. To assess the 360-degree synthesis capability, we randomly select 100 subjects from our Ours-360 dataset for testing, and the remaining for training.
@@
We evaluate the quality of our synthesized images using three quantitative metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). These measures enable us to effectively compare the synthetic images with the ground truth, ensuring a comprehensive assessment of the reconstruction quality. To evaluate identity similarity (CSIM) of the reconstructed 3D assets, we compute the cosine distance of face recognition features, building on the work of [20]. For assessing pose fidelity, we employ the Average Pose Distance (APD) [21].
Table 2: Running time for 3D head avatar generation from a single input image and rendering a 3D representation. PH-PTI denotes PanoHead-PTI, and SH-PTI denotes SphereHead-PTI.
Time H.NeRF EG3D-PTI PH-PTI SH-PTI Ours
Reconstruction 60s 120s 89.4s 96.2s 0.11s
Render 58ms 24ms 10.29ms 19.68ms 1.4ms
Table 3: Quantitative evaluation on the 100-person testset.
@@
4.2 Main Results
Qualitative Results.
Fig. 8 and Fig. 9 illustrate the novel view synthesis results after reconstruction from various methods, highlighting the superior performance of PanoLAM. Unlike prior approaches, PanoLAM enhances texture detail, maintains identity fidelity, and ensures multiview consistency. Due to the subtlety of multiview consistency in static images, we encourage viewing our supplementary video for a clearer demonstration. It is noteworthy that PanoLAM neither employs time-intensive test-time optimization nor relies on precise camera poses for accurate reconstruction, showcasing the efficacy of our pose-free framework. We achieve these results with 800X speedup.
Quantitative Results.
We benchmark our framework and baselines on the test set. For each subject, we evaluate 4 viewpoints: the front, back, left, and right. For the baseline method, we use their default configuration. As presented in Table 3, Our PanoLAM demonstrates exceptional reconstruction quality according to PSNR, SSIM, and LPIPS metrics. Furthermore, CSIM metrics indicate strong identity consistency of our methods. Remarkably, these achievements are coupled with fast reconstruction and rendering speeds, underscoring the effectiveness and efficiency of our approach.
@@ 4.2 Main Results
We analyze the feature aggregation mechanism from the spherical triplane of SphereHead to explore a better knowledge distillation strategy. In Fig. 6, we cast rays and visualize the aggregation weights of each sampled feature in the original ray marching. We plot 4 curves showing the weight values of 4 casting rays. As shown in the figure, multiple peaks occur near the geometry surface hit by each ray, indicating that we need to sample and aggregate multiple points on a ray to fetch valuable features from the spherical triplane fully. Moreover, each weight curve appears to be a different weight function, indicating that we cannot apply the same manual weight functions, e.g., a Gaussian distribution, for each ray. Thereafter, we propose to leverage learnable MLPs to learn the feature aggregation function from large-scale data. Experimental results in the ablation study validate such a finding and the effectiveness of our design.
@@ 4.3 Ablation Studies
Effect of Different Number of Gaussian Points.
Fig. 4 ablates the effect of different numbers of Gaussian points. As is shown in the figure, less number of Gaussian points are not able to reconstruct details like hair-strands and mustaches well, leading to blurry results in these regions.
@@
In Fig. 3, we visualize the effect of our coarse-to-fine Gaussian generation strategy. As is shown in the figure, the sparse points are easier to optimize and can be easily deformed to carve the shape of the person. However, a small number of Gaussian points cannot model texture details like hair strands. By leveraging the topology prior in the FLAME mesh, we effectively densify the points to learn more detailed textures. In contrast, directly predicting and regressing the dense point cloud results in insufficient optimization of Gaussian points. Many points are not well optimized for deformation and remain on the FLAME surface, which are invisible and wasted, causing dual-layer geometry in the hair region as well. Moreover, our coarse-to-fine strategy saves computation and GPU memory consumption since fewer points are needed for stacked cross-attention modules, which is shown in Table 4.
@@
Previous 3D head datasets lack subject diversity, as shown in Table 1. We ablate the effect of different numbers of subjects in Fig. 5. We keep all other configures the same and train our model with different numbers of subjects. We infer all models with the same testing subjects. The model trained on a limited number of subjects can not generalize well to unseen subjects, showing the effectiveness of our proposed large-scale, diverse dataset for network generalizability.
@@
In Fig. 7, we show the results of different query strategies from the spherical triplane. All strategies use the same feature renderer and frozen decoder from SphereHead for target image decoding. As shown in the figure, directly using the densified point cloud to query a single layer of features from the spherical triplane, as in [96], cannot fetch all valuable features, and the rendered images contain artifacts due to missing information. We then add 4 virtual cameras and cast rays from the camera center to each point and sample 32 layers of features near the point surface for feature aggregation. We set the margin between each sampled point to be the same as the fine ray marching stage in SphereHead. We try to assign manual weights sampled from a normal distribution centered at the selected point with standard deviation 1 and 10, and aggregate sampled features with a weighted sum. As is shown in the middle part of Fig. 7, the results look better but are still blurry. Our final solution utilizes learnable MLPs to learn the aggregation function for the NeRF-based rendering. As shown in the figure, ours gets the best results, showing the effectiveness of the proposed aggregation strategy.
Table 4: Ablation on our coarse-to-fine generation strategy on different numbers of Gaussian points. OOM denotes out of memory for an 80G GPU, and - denotes no results.
Metrics 5K D-30K C2F-30K D-100K C2F-100K
[…]

### v2

@@ 3.4 Large-scale Training Data Generation
BP4D-S 41 front PanoLAM-Front 48,135 $72^{\circ}$
Nersemble 222 front PanoLAM-360 117,186 $360^{\circ}$
@@
The scale of the training dataset is a key factor in enabling the generalizability of a trained model. However, obtaining a large dataset with diverse head assets is challenging. Though various datasets  (Yin et al., 2006; Zhang et al., 2014b; Hu et al., 2007; Savran et al., 2008; Cao et al., 2014; Cheng et al., 2018; Zhang et al., 2014a; Dai et al., 2020; Kirschstein et al., 2023; Cosker et al., 2011; Zhu et al., 2023; Wang et al., 2022; Pan et al., 2023; He et al., 2024; Martinez et al., 2024) have been proposed for 3D head modeling, as shown in Table 1, they have limited diversity of subjects, which hinders the generalization capability of learning-based models. Inspired by previous methods (Peebles et al., 2022; Cai et al., 2024; He et al., 2022; Wood et al., 2021) in various fields that model trained from synthesis data can generalize to real-world scenarios, we leverage pretrained priors from 3D GANs to generate a large-scale, diverse dataset for training. Specifically, we leverage two prior models for data generation, EG3D for multiview front face images synthesis, and SphereHead for 360 view images generation. Our observation is that EG3D generates a different distribution of images compared to SphereHead (e.g., diverse hairstyles and caps appeared in samples), while SphereHead can generate 360 images. The combination of two model spaces enhances the diversity of the generated images and improves the generalizability of learning-based models. However, bad cases occur in the generated dataset, hurting the model training. Therefore, we remove bad cases from the sampled images manually, ending up with 48,135 subjects in PanoLAM-Front and 117,186 subjects in PanoLAM-$360^{\circ}$ datasets. More details of the dataset construction and labeling pipeline are in Sec E.
@@ 3.5 Optimization and Regularization
In the training phase, we randomly select $N_{v}$ different views of images of the same subject. We randomly choose one as the reference image for the Gaussian full-head reconstruction, while the others serve as novel view images for prediction. Note that both branches in our framework render images for supervision. We ensure the accuracy of the rendered RGB images from each branch by comparing them with the ground truth target images, utilizing a combination of $\mathcal{L}_{1}$ loss and perceptual loss for supervision:
@@
Additionally, for the sparse and dense Gaussian Splatting branch, we render the silhouette and supervise it using $\mathcal{L}_{1}$ loss, referred to as $\mathcal{L}_{mask}$. For the spherical triplane branch, we supervise the extracted feature with those in the pretrained SphereHead by $\mathcal{L}_{1}$ loss, denoted as $\mathcal{L}_{st}$. The total loss function is a weighted sum as:
$\mathcal{L}=\lambda_{3}\mathcal{L}_{rgb}^{G}+\lambda_{4}\mathcal{L}_{rgb}^{ST}+\lambda_{5}\mathcal{L}_{mask}+\lambda_{6}\mathcal{L}_{st},$ (8)
where $\mathcal{L}_{rgb}^{G}$ denotes the RGB image loss for sparse and dense Gaussian Splatting and $\mathcal{L}_{rgb}^{ST}$ for the auxiliary spherical triplane branch, both denoted in Formula (7).
@@ 4 Experiments
Figure 4: Visualization of reconstruction and novel view synthesis of different methods.
@@ 4.1 Experiments Setting
Our framework is implemented in PyTorch, with the DINOv2 image feature extraction backbone frozen. The Transformer comprises 8 layers of basic blocks, featuring 16 attention heads and $C_{t}=512$ feature dimensions. The extracted features are translated into Gaussian attributes through three multilayer perceptron (MLP) layers. The network is trained over 100 epochs using the Adam optimizer and a cosine warm-up scheduler. Hyperparameters are empirically set as $N_{v}=6$, $\lambda_{1}=\lambda_{2}=1.0$, $\lambda_{3}=1.0$, $\lambda_{4}=0.1$, $\lambda_{5}=1.0$, $\lambda_{6}=0.0001$, $\epsilon_{O_{C}}=0.15$, $\epsilon_{O_{D}}=0.056$, $M_{D}=10K$, $K_{m}=32$, $\lambda_{area}=2.0e-6$, $H_{T}=W_{T}=64$.
@@
We employ the large-scale dataset that we have generated for training. To assess the 360 synthesis capability, we randomly select 100 subjects from our PanoLAM-360 dataset for testing, and the remaining for training.
@@
We evaluate the quality of our synthesized images using three quantitative metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). To evaluate identity similarity (CSIM) of the reconstructed 3D assets, we compute the cosine distance of face recognition features (Deng et al., 2022). For assessing pose fidelity, we employ the Average Pose Distance (APD) (Deng et al., 2019).
4.2 Main Results
Qualitative Results.
Fig. 4 illustrates the novel view synthesis results after reconstruction from various methods on the test set and out-of-domain real-world images, highlighting the superior performance of PanoLAM. Unlike prior approaches, PanoLAM enhances texture detail, maintains identity fidelity, and ensures multiview consistency. Compared to PanoHead and SphreHead, there are no artifacts in the background as well. It is noteworthy that PanoLAM neither employs time-intensive test-time optimization (getting 800X speedup) nor relies on precise camera poses for accurate reconstruction, showcasing the efficacy of our pose-free framework.
Quantitative Results.
Table 2: Quantitative evaluation on the testset.
@@ 4.2 Main Results
We benchmark our framework and baselines on the test set. For each subject, we evaluate metrics on 4 viewpoints: the front, back, left, and right, except the CSIM metric only uses the front view. We use the default configurations for baselines. As in Table 2, PanoLAM demonstrates exceptional reconstruction quality according to PSNR, SSIM, and LPIPS metrics. The CSIM metrics also indicate strong identity consistency of our methods. Remarkably, these achievements are coupled with fast reconstruction and rendering speeds, underscoring the effectiveness and efficiency of our approach.
Table 3: Running time for 3D avatar generation and rendering.
Time PH-PTI SH-PTI LAM Ours
Recon.(s) 89.4 96.2 1.4 0.11
Render(ms) 10.29 19.68 3.6 3.6
Table 4: Effect of Coarse-to-fine generation. OOM: Out of memory (80GB).
Metrics D-30K C2F-30K D-100K C2F-100K
Infer.(s) $\downarrow$ 0.49 0.082 - 0.11
[…]
Table 5: Module Ablations.
[…]
Full 23.494 0.793
w/o C2F 22.312 0.731
w/o ST ray agg. 22.156 0.715
2K subj. 19.391 0.699
@@
Figure 5: Analysis of ray marching aggregation weights in the original Spherical Triplane in SphereHead.
We analyze the feature aggregation mechanism from the spherical triplane of SphereHead to explore a better knowledge distillation strategy. In Fig. 5, we cast rays and visualize the aggregation weights of each sampled feature in the original ray marching. We plot 4 curves showing the weight values of 4 casting rays. Two phenomena occur as in the figure. Firstly, multiple peaks occur near the geometry surface hit by each ray, indicating that we need to sample and aggregate multiple points on a ray to fetch valuable features from the spherical triplane fully. Secondly, each weight curve appears to be a different weight function, indicating that we cannot apply the same manual weight functions, e.g., a Gaussian distribution, for each ray. Thereafter, we propose to leverage learnable MLPs to learn the feature aggregation function from large-scale data. Experimental results in the ablation study validate such a finding and the effectiveness of our design.
@@ 4.3 Ablation Studies
      Input   W/o coarse-to-fine    Coarse    Fine (Ours)
Figure 6: Effect of coarse-to-fine generation.
Fig. 6 shows the effect of our coarse-to-fine Gaussian generation strategy. The sparse points are easier to optimize and can be easily deformed to carve the shape of the person, but they cannot model texture details like hair strands. By densifying the points with topology prior in the FLAME mesh, our coarse-to-fine model can learn more detailed textures. In contrast, directly regressing the dense point cloud results in insufficient optimization of Gaussian points. Many points are not well optimized for deformation and remain on the FLAME surface, which are invisible and wasted, causing dual-layer geometry in the hair region as well. The quantitative results in Tab. 5 also support the effectiveness; without our coarse-to-fine training scheme (denoted as w/o C2F in the table), the performance drops. Moreover, as is shown in Table 5, our coarse-to-fine strategy saves computation and GPU memory consumption since fewer points are needed for stacked cross-attention modules. The network forward time and training GPU consumption are also reduced with our coarse-to-fine training strategy.
       Input        2K Subj.      Input    165K Subj.
Figure 7: Comparison of our network trained on different numbers of subjects.
@@
Previous 3D head datasets lack subject diversity, as shown in Table 1. We ablate the effect of different numbers of subjects in Fig. 7 and Table 5 (2K subj.). The model trained on a limited number of subjects cannot generalize well to unseen subjects, causing performance drops visually and quantitatively. This shows the effectiveness of our proposed large-scale, diverse dataset for network generalizability.
@@
Figure 8: Comparisons on different spherical triplane feature query and aggregation strategies.
In Fig. 8, we show the results of different query strategies from the spherical triplane. All strategies use the same feature renderer and frozen decoder from SphereHead for target image decoding. As shown in the figure, directly using the densified point cloud to query a single layer of features from the spherical triplane, as in (Zou et al., 2024), cannot fetch all valuable features, and the rendered images contain artifacts due to missing information. We then add 4 virtual cameras and cast rays from the camera center to each point and sample 32 layers of features near the point surface for feature aggregation. We set the margin between each sampled point to be the same as the fine ray marching stage in SphereHead. We try to assign manual weights sampled from a normal distribution centered at the selected point with standard deviation 1 and 10, and aggregate sampled features with a weighted sum. As is shown in the middle part of Fig. 8, the results look better but are still blurry. Our final solution utilizes learnable MLPs to learn the aggregation function for the NeRF-based rendering. As shown in the figure, ours gets the best results, showing the effectiveness of the proposed aggregation strategy. In Table 5, we also ablate the model without our proposed ray-based feature aggregation from spherical triplane (denoted w/o ST ray agg. in the table). The quantitative results drops and prove the effectiveness of the proposed module.